OpenStudy (anonymous):

9x^2+12x-32=0... solve?

OpenStudy (anonymous):

x=?

OpenStudy (anonymous):

x=4/3 x=-8/3

OpenStudy (anonymous):

use quadratic formula: http://www.purplemath.com/modules/quadform.htm

OpenStudy (saifoo.khan):

4 -- 3 -8 --- 3